A man invests $15,000 in an account that pays 8.5% interest per year, compounded quarterly. Find the amount after 6 years.
Give the answer tl the nearest cent. $________
Answer #1

P- 15000 8- 8.57 – 0.085 componded overleply n=u t= years ht UXG A = P (1+ A = 15000 (1 + 0095 24 A- 15ooo C U.085 u A = 1500